OpenStudy (anonymous):

The vertex form of the equation of a parabola is x = (y - 4)2 + 27. What is the standard form of the equation? A. x = y2 + 8y + 27 B. x = 4y2 - 8y + 43 C. x = y2 - 8y + 43 D. x = y2 + y + 15

OpenStudy (anonymous):

@Directrix @robtobey @rebeccaxhawaii @FortyTheRapper

OpenStudy (anonymous):

@caozeyuan

OpenStudy (caozeyuan):

you can expand the (y-4)^2

OpenStudy (caozeyuan):

the cheating way of doing it is to note that the square term is y^2, so B is gone

OpenStudy (anonymous):

x=ay²+by+c

OpenStudy (caozeyuan):

the linear term is -8y so C

OpenStudy (anonymous):

can u explain

OpenStudy (anonymous):

The vertex form of the equation of a parabola is y = (x + 3)2 + 53. What is the standard form of the equation? A. y = 6x2 + 9x+ 62 B. y = x2 + 6x + 62 C. y = x2 + 3x + 53 D. y = x2 + 53x + 42